### v2

@@ 3.2 Hyper Selecting Net
In this formula, $\mathbf{V_{s}}^{i}$ represents the $i$-th element of $\mathbf{V_{s}}$. The matrix $\mathbf{W}\in\mathbb{R}^{N\times P}$, referred to as the Weights Box, comprises $\mathbf{W}_{i}\in\mathbb{R}^{P}$ as its $i$-th column. The hyperparameter $N$ influences the total number of parameters, with $P$ being the count of parameters required for one Transformer Block. With $\mathbf{w}$ determined, it is then applied as parameters for the Transformer Block:
@@
Here, $\mathbf{x}$ signifies the input, and $\mathbf{x}^{\prime}$ is the resulting output. As the Transformer Blocks [12] are based on convolution, $\mathbf{w}\in\mathbb{R}^{P}$ is reshaped into 4-dimensional tensors to serve as convolution kernels for the Transformer Blocks. To economize on parameters, Transformer Blocks at the same decoder level share one Weights Box, and each Transformer Block is independently equipped with its own FCNN.
3.3 Rethinking Weights Generation via a Selecting Perspective
If we consider the Weights Box as part of the FCNN, it may not be immediately clear why such a direct derivation of weights from a neural network works well. This seeming lack of intuitiveness is what we aim to address. Before delving into the explanation, we establish a foundational proposition:
@@ 3.3 Rethinking Weights Generation via a Selecting Perspective
With this proposition, the following formula becomes evident:
$\mathbf{x}*\mathbf{w}=\mathbf{x}*\left(\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}\right)=\sum_{i=1}^{N}\left(\mathbf{x}*(\mathbf{V_{s}}^{i}\mathbf{W}_{i})\right)=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\left(\mathbf{x}*\mathbf{W}_{i}\right),$ (7)
where the constraint $\sum_{i}^{N}\mathbf{V_{s}}^{i}=1$ holds due to the property of Softmax operation. In essence, the convolution with the generated weights is equivalent to a weighted sum of convolutions between the input $\mathbf{x}$ and each kernel $\mathbf{W}_{i}$ in the Weights Box. This process can be seen as a process of selecting convolution kernels. For instance, with a Weights Box $\mathbf{W}\in\mathbb{R}^{2\times P}$ that includes a low-pass $\mathbf{W}_{1}$ and a high-pass $\mathbf{W}_{2}$, the rainy input $\mathbf{x}_{1}$ would ideally have a larger $\mathbf{V_{s}}^{1}$ to filter out high-frequency noise, while a smaller $\mathbf{V_{s}}^{2}$ would help retain details. Conversely, for a hazy input $\mathbf{x}_{2}$, a larger $\mathbf{V_{s}}^{2}$ would be necessary to mitigate low-frequency haze. Although real-world degradations can be more complex, our Hyper Selecting Net can adaptively select the appropriate weights. This insight into the selection process helps us understand HAIR, namely, for each input, the Hyper Selecting Net produces a tailored Selecting Vector and adaptively chooses the most suitable convolution kernel. Given that core operations such as convolution and matrix multiplication follow the distributive law over addition, HAIR is versatile and can be integrated into various architectures like Transformer [46] and Mamba [47].
3.4 A Brief Analysis of Model Complexity
In Section 2.3, we claim that our Hypernetworks-based method can work better than conditional embedding-based methods like AirNet [20] and PromptIR [24]. This section will give a simple proof for this point. In the context of all-in-one image restoration, we aim to find a function $f:\mathbb{R}^{3HW}\xrightarrow{}\mathbb{R}^{3HW}$ to map a degraded image $\mathbf{X}$ to a restored image $\hat{\mathbf{X}}$ by minimizing the distance between $\hat{\mathbf{X}}$ and the ground truth $\mathbf{Y}=y(\mathbf{X})$. [^1] For the sake of discussion, the following tensors are generally regarded as flattened one-dimensional vectors. Mainstream methods typically utilize a neural network $g(\mathbf{X},e(\mathbf{X}))$ to learn the mapping, where $e(\mathbf{X})\in\mathbb{R}^{k}$ stands for the degradation information (degradation embedding) from the input image, such as the prompt in PromptIR [24]. These methods directly send $e(\mathbf{X})$ together with $\mathbf{X}$ into the network to get $\hat{\mathbf{X}}$. For our HAIR, the neural network can be formulated as $h(\mathbf{X};\theta(e(\mathbf{X})))$, where $\theta$ is a function that maps ${e(\mathbf{X})}$ into the parameters of $h$. We define the distance between the two functions as:
@@ 3.4 A Brief Analysis of Model Complexity
Given that vector functions can be complex, we define a scalar function $f^{i}:\mathbb{R}^{M}\xrightarrow{}\mathbb{R},i=1,2,3\cdots M.$ of a vector function $f:\mathbb{R}^{M}\xrightarrow{}\mathbb{R}^{M}$. $i$ means $f^{i}$ produce the $i$-th value of $f$, then we have:
$=\min_{g}\max_{i}\max_{\mathbf{X}}|g^{i}(\mathbf{X})-y^{i}(\mathbf{X})|$
$=\min_{g}\max_{\mathbf{X}}|g^{k(g)}(\mathbf{X})-y^{k(g)}(\mathbf{X})|.$
Since the value range of $i$ is limited, given the function $g$, we can always find an integer $k(g)$ to replace $\max_{i}$. To complete the proof, we need some assumptions.
Assumption 1.
The target function $y^{i}\in\mathcal{W}_{r,3HW},i=1,2,\cdots 3HW$. The Sobolev space $\mathcal{W}_{r,3HW}$ is the set of functions that are $r$-times differentiable with all $r$-th order derivatives being continuous and bounded by the Sobolev norm $\leq 1$, defined on $\mathbb{R}^{3HW}$.
Intuitively, since $y^{i}$ maps the degraded image to a single pixel of the clean image, a little disturbance on the degraded image should only bring a little difference to the single pixel, so $y^{i}$ should be smooth and differentiable or at least continue. What’s more, the domain and range of $y$ are restricted to $[0,1]^{3HW}$, with most of the pixels far less than 1, the Sobolev norm of $y_{i}$ can be generally less than 1. So this assumption generally holds.
Assumption 2.
For various functions (especially neural networks) $g$, the integer $k(g)$ in Eq. 9 remains the same.
In our context, the various $g$ represents the same network with different weights during training. No matter how the parameters update, the most "difficult" degraded image generally remains the same one, e.g., the one with very severe degradation. Within this "difficult" image, the most "difficult" pixel should also remain the same, e.g., the pixel that varies dramatically from its clean version. In this way, we simplify the vector function $g$ to scalar function $g^{k(g)}$. With these two assumptions, as well as incorporating Theorem 2-4 and all the assumptions in [32], we can obtain the following conclusion.
@@
For conditional embedding-based all-in-one image restoration methods $g(\mathbf{X},e(\mathbf{X}))$ like PromptIR [24], to achieve error $d(g,y)\leq\epsilon$, the complexity (number of parameters) of the corresponding model is at least $N_{g}=\Omega(\epsilon^{-3HW-k})$.
@@
For Hypernetworks based methods $h(\mathbf{X};\theta(e(\mathbf{X})))$ like our HAIR, to achieve error $d(h,y)\leq\epsilon$, the complexity of the corresponding model is $N_{\theta}=O(\epsilon^{-3HW/r})$.
Please refer to paper [32] and Section 6.1 for detailed premises and proof of Theorem 1 and 2. The two Theorems demonstrate that to achieve the same error $\epsilon$, our proposed HAIR requires fewer parameters than conditional embedding-based methods. (As long as $\epsilon$ is small enough.)
@@ 4 Experiments
We conduct experiments by rigorously adhering to the methodologies established in prior research on general image restoration [25, 24], under two distinct experimental configurations: (a) All-in-One and (b) Single-task. In the All-in-One configuration, a singular model is trained to handle multiple types of degradation, encompassing both three and five unique degradations. In contrast, the Single-task configuration involves the training of individual models dedicated to specific restoration tasks.
@@ 4.1 Experimenting Preparation
Implementation Details. Our Res-HAIR framework is designed to be end-to-end trainable, eliminating the need for pre-training any of its components. The architecture comprises a 4-level encoder-decoder structure, with each level equipped with a varying number of Transformer blocks. Specifically, the number of blocks increases progressively from level-1 to level-4, following the sequence [4, 6, 6, 8]. HyperTrans Blocks are employed throughout level-4 and the decoding stages of levels 1-3. Additionally, the Weights Box parameter $N$ is set according to the sequence [5, 7, 7, 9] for each respective level. In the all-in-one setting, the model is trained with a batch size of 32, while in the single-task setting, a batch size of 8 is used. The optimization of the network is guided by an L1 loss function, employing the AdamW optimizer [48] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is set to $42e-$ 4 for the initial 150 epochs, and we change it to $52e-$ 5 for the final 10 epochs. To enhance the training data, input patches of size 128 $\times$ 128 are utilized, with random horizontal and vertical flips applied to the images to augment the dataset.
Datasets. We follow the approaches in previous studies [20, 24, 25] for our All-in-One and single-task experiments, using the following datasets. For single-task image denoising, we use the BSD400 [49] and WED [50] datasets, adding Gaussian noise at levels $\sigma\in\{15,25,50\}$ to generate training images. Testing is performed on the BSD68 [51] and Urban100 [52] datasets. For deraining, we employ the Rain100L dataset [53]. Dehazing experiments utilize the SOTS dataset [54]. Deblurring and low-light enhancement tasks use the GoPro [55] and LOL-v1 [56] datasets respectively. To create a comprehensive model for all tasks, we combine these datasets and train them in a setting that covers three or five types of degradations. For single-task models, training is conducted for the same duration on the respective dataset.
Table 1: Comparison to state-of-the-art on three degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best and second best performances are highlighted. Our method sets a new state-of-the-art on average across all benchmarks while being significantly more efficient than previous work. ‘-’ represents unreported results.
@@ 4.2 Results
All-in-One: Three degradations. We conducted a comparative evaluation of our All-in-One image restoration framework against several state-of-the-art specialized methods, including BRDNet [57], LPNet [58], FDGAN [59], DL [23], MPRNet [21], AirNet [20], PromptIR [24], and DaAIR [26]. Each of these methods was trained to handle the tasks of dehazing, deraining, and denoising. As illustrated in Table 1, our approach excels as the leading contender, offering a significant average improvement of $0.64$ dB over PromptIR [24]. Our method notably outperforms existing benchmarks on BSD68 and Rain100L datasets, setting a new precedent by exceeding the performance of the previous best methods by margins of $0.19$ dB and $1.49$ dB respectively.
All-in-One: Five Degradations. Following recent studies [25, 26], we have extended our approach to include deblurring and low-light image enhancement, thereby broadening our three-degradation framework to a more complex All-in-One setting. As demonstrated in Table 2, our method excels by learning specialized models for each degradation type while effectively capturing the shared characteristics between tasks. It achieves an average improvement of $0.6$ dB over PromptIR [24] and $0.13$ dB over DaAIR [26], establishing a new benchmark for state-of-the-art (SOTA) performance across all five benchmarks. Additionally, we compared our method with general image restoration models trained in the same All-in-One setting. Notably, our method significantly outperforms our baseline (i.e. Restormer [12]) in various tasks. This comparison highlights the strong capability of our method as a versatile plug-in-and-play module, enhancing the performance of existing restoration architectures with minimal integration complexity.
Table 2: Comparison to state-of-the-art on five degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images with $(^{\ast})$ denoting general image restorers, others are specialized all-in-one approaches. Best and second best performances are highlighted.
@@
Single-Degradation Results. To evaluate the effectiveness of our proposed method, we provide results in Table 3, which show the performance of individual instances of our method trained under a single degradation protocol. Specifically, the single-task variant dedicated to deraining consistently achieves higher performance than PromptIR [24] and DaAIR by margins of $1.96$ dB and $1.22$ dB, respectively. When applied to image denoising, our method also demonstrates superiority over the aforementioned approaches, with an improvement of $0.21$ dB and $0.47$ dB, respectively. These results underscore the significant enhancements in performance delivered by our method.
Table 3: Comparison to state-of-the-art for single degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best and second best performances are highlighted. Our method excels prior work on deraining and denoising.
@@
Visual Results The visual results captured under three degradation scenarios are presented in Fig. 4. These results clearly demonstrate the superiority of our method in terms of visual quality: In the denoising task at $\sigma=25$, Res-HAIR retains more fine details compared to other methods; in the deraining task, Res-HAIR effectively removes all rain streaks, whereas the compared methods exhibit noticeable remnants; in the dehazing task, Res-HAIR may not closely resemble the Ground Truth, yet it provides the most visually pleasing outcome and even clears the original haze present in the Ground Truth. Collectively, these observations confirm the efficacy of our approach in enhancing image quality across different degradation conditions.
Figure 4: Visual comparison of all-in-one methods in all-in-one degradation settings.

### v4

@@ 3.1.2 Hyper Selecting Net (HSN)
In this formula, $\mathbf{V_{s}}^{i}$ represents the $i$-th element of $\mathbf{V_{s}}$. The matrix $\mathbf{W}\in\mathbb{R}^{N\times P}$, referred to as the Weight Box, comprises $\mathbf{W}_{i}\in\mathbb{R}^{P}$ as its $i$-th row. The hyperparameter $N$ influences the total number of parameters, with $P$ being the count of parameters required for one corresponding module (i.e. one Transformer Block in this example). With $\mathbf{w}$ determined, it is then used as the parameters of the corresponding module. So the operation of the "HyperTrans Block" in Fig. 3 can be represented as:
@@
Here, $\mathbf{x}$ is the input and $\mathbf{x}^{\prime}$ is the output. As the Transformer Block (see Sec. 9) is based on convolution, $\mathbf{w}\in\mathbb{R}^{P}$ is reshaped into four-dimensional tensors to serve as convolution kernels for the Transformer Block. To reduce the number of parameters, the transformer blocks at the same decoder level share one weight box, and each transformer block is independently equipped with its own FCNN.
3.2 Proposed Res-HAIR Model
Integrating HAIR into an existing image-to-image network involves a simple two-step process: (1) Insert the DAC at the junction between the first and second halves of the network to produce a Global Information Vector (GIV) from the features of the first half. (2) Incorporate HSNs into the second half of the network, using the GIV to generate weights for all modules within this section dynamically. In this work, we integrate our proposed HAIR module into the popular image restoration model Restormer [60], and propose our All-in-One model Res-HAIR. Although we only use the example of integrating HAIR with Restormer to illustrate the process, this example is representative, and integration with other networks follows in a virtually identical way.
Overall Architecture. As shown in Fig. 3, our network architecture is consistent with Restormer [60]. Given a certain degraded image input $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$, Res-HAIR utilizes a $3\times 3$ convolution to transform $\mathbf{X}$ into feature embeddings $\mathbf{F_{0}}\in\mathbb{R}^{H\times W\times C}$, where $C$ denotes the number of channels. These feature embeddings are then proceeded through a 4-level hierarchical encoder-decoder, resulting in deep features $\mathbf{F_{r}}$. Each level of the encoder-decoder incorporates several Transformer blocks, with an increasing number from the top to the bottom level, ensuring computational efficiency. The left three blue "Transformer Blocks" function as an Encoder, designed to extract features from $\mathbf{F_{0}}$ and ultimately produce a global feature map $\mathbf{F_{g}}\in\mathbb{R}^{\frac{W}{8}\times\frac{H}{8}\times 8C}$ with a large receptive field. $\mathbf{F_{g}}$ is then input into the DAC to yield a Global Information Vector (GIV) $\mathbf{V_{g}}\in\mathbb{R}^{2C}$. The right four orange "HyperTrans Blocks" operate as a Decoder, aiming to adaptively fuse features at each Decoder level based on degradation, culminating in a restored image $\hat{\mathbf{X}}$. Specifically, at each Decoder level, the HyperTrans Block receives two inputs, i.e. the input feature and the GIV $\mathbf{V_{g}}$. $\mathbf{V_{g}}$ is fed into the Hyper Selecting Nets to generate weights for the corresponding Transformer Blocks, which are then applied to the input feature to produce the output. It is important to note that all HyperTrans Blocks use the same $\mathbf{V_{g}}$ derived from $\mathbf{F_{g}}$. Since the weights in the Decoder are generated based on $\mathbf{V_{g}}$, which contains the degradation information of $\mathbf{X}$, our method can thus adapt to different degraded images.
3.3 Theoretical Analysis
In this subsection, we first interpret the generated weights from a selection perspective and then provide a brief theoretical analysis of model complexity.
3.3.1 Hypernetworks from a Selection Perspective
Considering the Weight Box as part of the FCNN, the direct generation of weights may still appear somewhat opaque. To clarify this process, we first introduce a proposition before diving into the detailed explanation.
@@ 3.3.1 Hypernetworks from a Selection Perspective
The following formula then becomes evident:
$\mathbf{x}*\mathbf{w}=\mathbf{x}*\left(\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}\right)=$ (7)
$\sum_{i=1}^{N}\left(\mathbf{x}*(\mathbf{V_{s}}^{i}\mathbf{W}_{i})\right)=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\left(\mathbf{x}*\mathbf{W}_{i}\right),$
where $\sum_{i}^{N}\mathbf{V_{s}}^{i}=1$ holds due to the property of Softmax operation. Essentially, the convolution with the generated weights is equivalent to a weighted sum of convolutions between the input $\mathbf{x}$ and each kernel $\mathbf{W}_{i}$ in the Weight Box. This process can be seen as a process of selecting convolution kernels. If we view each $\mathbf{W}_{i}$ as an expert, it’s also like the Mixture-of-Experts pattern used in [56]. For instance, with a Weight Box $\mathbf{W}\in\mathbb{R}^{2\times P}$ that includes a low-pass $\mathbf{W}_{1}$ and a high-pass $\mathbf{W}_{2}$, the rainy input $\mathbf{x}_{1}$ would ideally have a larger $\mathbf{V_{s}}^{1}$ to filter out high-frequency noise, while a smaller $\mathbf{V_{s}}^{2}$ would help retain details. In contrast, for a hazy input $\mathbf{x}_{2}$, a larger $\mathbf{V_{s}}^{2}$ would be necessary to mitigate low-frequency haze. Although real-world degradations can be more complex, our HSN can adaptively select the appropriate weights. This insight into the selection process helps us understand HAIR, i.e., HSN produces a tailored Selecting Vector and adaptively chooses the most suitable convolution kernel for each input. Given that core operations such as convolution and matrix multiplication follow the distributive law over addition, HAIR is universal and can be integrated into architectures like Transformer [49] and Mamba [18].
3.3.2 A Brief Analysis of Model Complexity
In Sec. 2.2, we claim that our Hypernetworks-based method can be more efficient than conditional embedding-based methods such as AirNet [27] and PromptIR [38]. This section will provide a simple proof for this point. In the context of All-in-One image restoration, we aim to find a function $f:\mathbb{R}^{3HW}\xrightarrow{}\mathbb{R}^{3HW}$ to map a degraded image $\mathbf{X}$ to a restored image $\hat{\mathbf{X}}$ by minimizing the distance between $\hat{\mathbf{X}}$ and the ground truth $\mathbf{Y}=y(\mathbf{X})$ [^1] For the sake of discussion, the following tensors are generally regarded as flattened one-dimensional vectors.. Mainstream methods typically utilize a network $g(\mathbf{X},e(\mathbf{X}))$ to learn the mapping, where $e(\mathbf{X})\in\mathbb{R}^{k}$ contains the degradation embedding of the input image, such as the prompt in PromptIR [38]. These methods send $e(\mathbf{X})$ together with $\mathbf{X}$ into the network to get $\hat{\mathbf{X}}$. For our HAIR, the network can be formulated as $h(\mathbf{X};\theta(e(\mathbf{X})))$, where $\theta$ is a function that maps ${e(\mathbf{X})}$ into the parameters of $h$. We define the distance between functions as
@@ 3.3.2 A Brief Analysis of Model Complexity
Then we have the following 2 theorems.
@@
For conditional embedding-based All-in-One image restoration methods $g(\mathbf{X},e(\mathbf{X}))$ like PromptIR [38], to achieve error $d(g,y)\leq\epsilon$, the complexity (number of parameters) of the corresponding model is at least $N_{g}=\Omega\left(\epsilon^{-\min(3HW+k,6HW)}\right)$, where $k$ is the dimensionality of the embedding vector. Typically $k=\Omega(HW)$, $k\geq 3HW$.
@@
For Hypernetworks based methods $h(\mathbf{X};\theta(e(\mathbf{X})))$ like our HAIR, to achieve error $d(h,y)\leq\epsilon$, the complexity of the corresponding model is $N_{\theta}=O(\epsilon^{-3HW/r})$, where $r\geq 1$.
The two theorems demonstrate that to achieve the same error $\epsilon$, our proposed HAIR requires fewer parameters than conditional embedding-based methods as long as $\epsilon$ is small enough. Please refer to [15] and Sec. 6 in supplementary materials for detailed proof of Theorem 1 and 2.
@@ 4 Experiments
The experimental settings follow previous research on general image restoration [62, 38] in two different configurations: (a) All-in-One and (b) Single task. In the All-in-One configuration, a singular model is trained to handle multiple types of degradation, encompassing both three and five unique degradations. In contrast, the Single-task configuration involves training individual models dedicated to specific restoration tasks.
@@ 4.1 Experimenting Preparation
Implementation Details. Our Res-HAIR method is designed to be end-to-end trainable, eliminating the need for pre-training any of its components. The architecture comprises a 4-level encoder-decoder structure, with each level equipped with a varying number of Transformer blocks. Specifically, the number of blocks increases progressively from level-1 to level-4, following the sequence [4, 6, 6, 8]. HyperTrans Blocks are employed throughout level-4 and the decoding stages of levels 1-3. Additionally, the Weight Box parameter $N$ is set according to the sequence [5, 7, 7, 9] for each respective level. In the All-in-One setting, the model is trained with a batch size of 32, while in the single-task setting, a batch size of 8 is used. The network optimization is guided by an L1 loss function, employing the AdamW optimizer [31] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is set to $42e-$ 4 for the initial 150 epochs and then changed to $52e-$ 5 for the final 10 epochs. To enhance the training data, input patches of size 128 $\times$ 128 are utilized, with random horizontal and vertical flips applied to the images to augment the dataset.
Datasets. We follow previous approaches [27, 38, 62] for our All-in-One and single-task experiments, using these benchmark datasets. For single-task image denoising, we use the BSD400 [3] and WED [33] datasets, adding Gaussian noise with levels $\sigma\in\{15,25,50\}$ to generate training images. Testing is performed on the BSD68 [34] and Urban100 [21] datasets. For deraining, we employ the Rain100L dataset [53]. Dehazing experiments utilize the SOTS dataset [26]. Deblurring and low-light enhancement tasks use the GoPro [36] and LOL-v1 [51] datasets, respectively. To create a comprehensive model for all tasks, we combine these datasets and train them in a setting that covers three or five types of degradations. For single-task models, training is conducted on the respective dataset.
Table 1: Comparison to state-of-the-art on three degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images with $(^{\ast})$ denoting methods that are not blind (need prior information of the degradation). Best and second best performances are highlighted.
@@ 4.2 Results
All-in-One: (1) Three degradations. We conducted a comparative evaluation of our All-in-One image restoration model against several state-of-the-art specialized methods, including BRDNet [46], LPNet [16], FDGAN [12], DL [14], MPRNet [59], AirNet [27], PromptIR [38], InstructIR [9] and DaAIR [56]. Each of these methods was trained to handle the tasks of dehazing, deraining, and denoising. As illustrated in Tab. 1, our approach can significantly outperform other competing methods, e.g. an average improvement of $0.64$ dB over PromptIR. Our method notably outperforms existing benchmarks on SOTS and Rain100L datasets, by exceeding the performance of the previous best methods by margins of $0.4$ dB and $0.61$ dB respectively. (2) Five Degradations. Following recent studies [62, 9], we have extended our approach to deblurring and low-light image enhancement, thus extending the three-degradation setting to a more complex five-degradation setting. As demonstrated in Tab. 2, our method excels by learning specialized models for each degradation type while effectively capturing the shared characteristics between tasks. It achieves an average improvement of $2.03$ dB over PromptIR and $0.82$ dB over the non-blind method InstructIR, establishing a new benchmark for state-of-the-art performance across all five benchmarks. Notably, our method significantly outperforms our baseline (i.e. Restormer) in various tasks. This comparison highlights the strong capability of our method as a versatile plug-in-and-play module, enhancing the performance of the existing model with a small amount of integration complexity, e.g. only adding 3M parameters compared with Restormer.
Table 2: Comparison to state-of-the-art on five degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images with $(^{\ast})$ denoting methods that are not blind (need prior information of the degradation type). Best and second best performances are highlighted.
@@
Single-Degradation Results. To evaluate the effectiveness of our proposed method, we provide results in Tab. 3, which show the performance of individual instances of our method trained under a single degradation protocol. Specifically, the single-task variant dedicated to deraining consistently achieves higher performance than PromptIR and InstructIR by margins of $1.96$ dB and $1.02$ dB, respectively. When applied to image denoising, our method also demonstrates superiority over the aforementioned approaches, with an average improvement of $0.21$ dB and $0.47$ dB, respectively. These results underscore the significant performance improvements delivered by our method.
Table 3: Comparison to state-of-the-art for single degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best and second best performances are highlighted. Denoising results are from Urban100  [21].
@@
Visual Results. The visual results captured under three degradation scenarios are presented in Fig. 4. These results clearly demonstrate the superiority of our method in terms of visual quality. In the denoising task at $\sigma=25$, Res-HAIR retains more fine details compared to other methods; in the deraining task, Res-HAIR effectively removes all rain streaks, whereas the compared methods contain obvious rain streaks; in the dehazing task, Res-HAIR may not closely resemble the ground-truth, yet it provides the most visually pleasing outcome and even clears the original haze present in the ground-truth. Overall, these visual observations confirm the efficacy of our approach in enhancing image quality across different degradation conditions.
Figure 4: Visual comparison on various degradation settings. Zoom for better visual effects.